If excluded volume is taken zero compressibility factor z is
a)1-a/RTV
b)1+Pb/RT
c)PV/RT
d)RT/PV

Grade:11

3 Answers

Himanshu Dogra
13 years ago

If excluded volume is taken zero, Vander Wall's Equation becomes:

(P +  an2

(P +  a

 

 

 

 

 

 

 

(P +  an2

Himanshu Dogra
13 years ago

Vander Wall's Equation (for 0 excluded volume) for 1 mole of an ideal gas:

(P + a/V

 

 

 

 

 

 

vikas askiitian expert
13 years ago

vander valls equation is

(p+a/v2)(v-b) = RT

if excluded volume is 0 then correction term in volume means 'b' is 0 ...

now , eq becomes

(p + a/v2)v = RT

pv + a/v = RT

pv/RT + a/vRT = 1

z + a/vRT = 1

z = 1 - a/vRT

option a is correct